if a train is moving at the rate of 1 kilometer in 1 minute and 20 seconds how far will it travel in 1 hour
Answer by edjones(7311)

Do a proportion.
km/t=km/t
1/(4/3)=x/60
(4/3)means 4/3 of a minute
multiply 60 times both sides. x=60/(4/3)
x=60*3/4=45 km
EdJones